Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
menzione - 22:58 Friday 23 February 2024 (63389) Print this report
Operator Report - Afternoon shift

Upon arrival I found ITF relocking. It reached the lock easily LN3.
The planned activity on TCS - ETM RH tuning (Nardecchia) started at 16:00 UTC and went on without particular problems till the unlock at 19:12 UTC.
After the unlock I had to rearm vbias of photodiode SIB2_B2_PD1, SIB2_B2_PD3.
During the usual cross-alignment at ACQUIRE_DRMI the central spot moved in a unusual way. I had to engaged the fixed setpoint from SDB1_LC.
Unfortunately we lost time becouse the Sat Server crashed during the realignment.
Upon solved those issues the lock acquisition restarted but the bad weather conditions prevented a stable lock.

ITF lef tin AUTORELOCK_FAILSAFE

Images attached to this report
Search Help
×

Warning

×